RC Meeting: September 6th, 2020

From clas12-run
Jump to: navigation, search

Back to Run Group F | Daily Run Group F Coordination Meetings


2020

March
19, 20, 21, 22, 23, 24
July
31
August
1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31
September
1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20 21

RC:

Summary of Sep. 6th

Owl shift

  1. Shift summary by the shift expert: https://logbooks.jlab.org/entry/3841968
  2. Issues: Ratty beam, DAQ issues.
  3. Recorded data: 13.3 M events on H2

Day shift

  1. Shift summary by the shift expert: https://logbooks.jlab.org/entry/3842229
  2. Issues: Still ratty beam.
  3. Recorded data: 26.7 M events on H2

Swing shift

  1. Shift summary by the shift expert: https://logbooks.jlab.org/entry/3842510
  2. Issues: Switched to empty (1-3 psig H2) target and increased beam current to 300 nA. Maybe a mistake - the 2H01 vacuum problem came roaring back with a vengeance. Stopped beam and explored in the Hall. "Fixed itself" for now - continue with 4He. Followed by Moller measurement (80.6%) and back to D2.
  3. Recorded data: 5.4 M events on MT, 6.8 M events on 4He.

Whole day summary

  1. Total recorded data today: 63 M events (H2/MT/He/D2)
  2. Part-II total accumulated data on D2: 1.57 B events
  3. Concerns: Continued beam troubles and equipment failures in accelerator; 2H01 vacuum issue (may have to fix beam exit window).


Tomorrow daily RC meeting will be at usual time 12:00 PM using the link https://bluejeans.com/7576835804/

Recorded data numbers

Number of events recorded